# Balanced Multi-view Clustering

### v1

@@ Abstract
Multi-view clustering (MvC) aims to integrate information from different views to enhance the capability of the model in capturing the underlying data structures. The widely used joint training paradigm in MvC is potentially not fully leverage the multi-view information, since the imbalanced and under-optimized view-specific features caused by the uniform learning objective for all views. For instance, particular views with more discriminative information could dominate the learning process in the joint training paradigm, leading to other views being under-optimized. To alleviate this issue, we first analyze the imbalanced phenomenon in the joint-training paradigm of multi-view clustering from the perspective of gradient descent for each view-specific feature extractor. Then, we propose a novel balanced multi-view clustering (BMvC) method, which introduces a view-specific contrastive regularization (VCR) to modulate the optimization of each view. Concretely, VCR preserves the sample similarities captured from the joint features and view-specific ones into the clustering distributions corresponding to view-specific features to enhance the learning process of view-specific feature extractors. Additionally, a theoretical analysis is provided to illustrate that VCR adaptively modulates the magnitudes of gradients for updating the parameters of view-specific feature extractors to achieve a balanced multi-view learning procedure. In such a manner, BMvC achieves a better trade-off between the exploitation of view-specific patterns and the exploration of view-invariance patterns to fully learn the multi-view information for the clustering task. Finally, a set of experiments are conducted to verify the superiority of the proposed method compared with state-of-the-art approaches both on eight benchmark MvC datasets and two spatially resolved transcriptomics datasets. The demo code of this work is publicly available at https://github.com/guanyuezhen/BMvC.
@@ 1 Introduction
  • We perform extensive experiments on eight benchmark multi-view datasets to verify the effectiveness of the proposed method. Additionally, our BMvC method is extended to apply to the spatially resolved transcriptomics data clustering task to demonstrate the efficacy of the proposed method.
@@ 4.3.3 Parameter Sensitivity Analysis
4.4 Application for Spatially Resolved Transcriptomics Data Clustering
Figure 4: Illustration of proposed BMvC method on the spatially resolved transcriptomics data clustering task.
The spatial arrangement of various cell types is closely linked to the functionality of complex tissues. Recent advancements in spatial transcriptomics (ST) have transformed our capacity to investigate this relationship. By combining transcriptomic data with their respective spatial contexts, researchers can achieve a more profound understanding of the biological roles of these tissues. Furthermore, spatial information aids in mapping intercellular communication, thereby deepening our comprehension of cell interactions and the regulatory mechanisms that influence cell behavior. The fusion of spatial and transcriptomic data provides a more holistic view of complex tissue functions. However, the heterogeneity between gene expression and spot spatial location results in an imbalanced multi-view learning issue in previous spatial transcriptomics clustering methods. For example, the spatial continuity of cellular tissue makes it preferable to learn the information from the spot spatial location rather than the gene expression for clustering. Both the gene expression and spot spatial location information are important for revealing the cellular relationships. Thus, it is necessary to formulate a balanced multi-view learning framework for the spatially resolved transcriptomics data clustering task.
In this section, we transfer our proposed BMvC method to deal with the spatially resolved transcriptomics data clustering task, as shown in Fig. 4. Specifically, two graph convolutional neural networks are employed to extract the transcriptomic expression-related representation and spatial distribution-related representation to comprehensively depict each cell. Then, the feature concatenation operator consequently with a linear layer is utilized to fusion such two different types of information. As done in previous methods [70, 71], the ZINB decoder [72], which captures the complex global information of data, is introduced to reconstruct the gene expression. Finally, the view-specific contrastive regularization is leveraged to achieve more balanced multi-view learning in the spatially resolved transcriptomics data clustering task.
4.4.1 Datasets
We evaluate the proposed method on two spatially resolved transcriptomics data:
Human Breast Cancer (HBC[73]: This dataset features 20 meticulously annotated regions categorized into four main morphological types.
Mouse Brain Anterior Tissue (MBA[74]: This dataset is labeled with 52 distinct class areas and captured using the 10X Visium device, and shows minimal inter-class variation.
4.4.2 Compared Methods
To verify the superiority of proposed method, we compare it with twelve benchmark methods: K-means, SCANPY [75], SpaGCN [76], DeepST [77], SCGDL [78], GraphST [79], Spatial-MGCN [80], stLearn [81], CellPLM [82], stMMR [83], MAFN [70], and stGCL [71]. In our experiments, we set the parameters of all compared methods as suggestions according to their papers to implement their best clustering performance.
[…]
Datasets HBC MBA
Methods NMI ARI NMI ARI
K-means 51.45 35.12 58.75 24.45
SCANPY[G. Biol.’18][75] 51.59 49.15 45.29 24.43
SpaGCN[Nat. Methods’21][76] 64.34 56.12 62.72 34.23
DeepST[NAS’22][77] 68.75 53.13 56.71 25.63
SCGDL[BIB’23][78] 43.81 35.32 64.30 26.95
GraphST[Nat. Com.’23][79] 67.00 52.65 68.89 38.59
Spatial-MGCN[BIB’23][80] 69.81 64.32 71.13 42.46
stLearn[bioRxiv’20] [81] 62.64 55.26 65.52 38.42
CellPLM[ICLR’24][82] 33.13 19.12 57.17 24.77
[…]
MAFN[TKDE’24][70] 69.32 61.88 72.76 43.58
stGCL[ACM MM’24][71] 69.27 65.62 69.25 47.63
[…]
4.4.3 Spatial Clustering Performance
We report the spatial clustering performance in terms of ARI and NMI of all compared methods in Tab. III. From the results, we note that our BMvC method achieves superior results compared with previous well-designed spatial clustering methods on the two datasets. For example, our BMvC method obtains the best performance in terms of ARI on HBC and MBA datasets, and is better than the second performer by about 1.36 and 2.81 percentages. Previous spatial clustering methods usually follow the joint-training paradigm to perform multi-view learning. As mentioned in Section 3.2.2, the joint-training paradigm poses an imbalanced learning problem, making the view-specific encoders under-optimized. Thus, such joint-training paradigm based methods have limited spatial clustering performance. In contrast, our BMvC method introduces a view-specific contrastive regularization to find a better trade-off between the exploitation of view-specific patterns and the exploration of view-invariance patterns to fully learn the multi-view information for the clustering task.
Figure 5: Clustering distribution visualization results of compared methods on HBA dataset.
Figure 6: Clustering distribution visualization results of compared methods on MBA dataset.
4.4.4 Visualization Analysis
The clustering distribution visualization results of all compared methods are presented in Figs. 5 and 6. Compared with other methods, our method provides more compact and region-consistent visualization results, especially in the left bottom corner and left part of the HBC and MBA datasets, respectively. Furthermore, the proposed method obtains clearer spatial boundaries, indicating more accurate cell clustering. Such visualization superiority of the proposed method remarkably demonstrates the effectiveness of the proposed method and the efficiency of the well-designed view-specific contrastive regularization in dealing with the imbalanced multi-view problem.
@@ 5 Conclusions
In this paper, we proposed a novel balanced multi-view clustering method to achieve more balanced multi-view learning and further improve the clustering performance. We first analyzed the imbalanced multi-view clustering problem existing in the joint-training paradigm from the gradient view. Then, we introduced a view-specific contrastive regularization to make a better trade-off between the exploitation of view-specific patterns and the exploration of view-invariance patterns to fully learn the multi-view information for the clustering task. Additionally, the theoretical analysis was provided to verify the effectiveness of the VCR from the gradient perspective. Finally, Extensive experiments on various benchmark multi-view clustering datasets and spatially resolved transcriptomics datasets were conducted to verify the efficacy of our method.

### v3

@@ Abstract
Multi-view clustering (MvC) aims to integrate information from different views to enhance the capability of the model in capturing the underlying data structures. The widely used joint training paradigm in MvC is potentially not fully leverage the multi-view information, since the imbalanced and under-optimized view-specific features caused by the uniform learning objective for all views. For instance, particular views with more discriminative information could dominate the learning process in the joint training paradigm, leading to other views being under-optimized. To alleviate this issue, we first analyze the imbalanced phenomenon in the joint-training paradigm of multi-view clustering from the perspective of gradient descent for each view-specific feature extractor. Then, we propose a novel balanced multi-view clustering (BMvC) method, which introduces a view-specific contrastive regularization (VCR) to modulate the optimization of each view. Concretely, VCR preserves the sample similarities captured from the joint features and view-specific ones into the clustering distributions corresponding to view-specific features to enhance the learning process of view-specific feature extractors. Additionally, a theoretical analysis is provided to illustrate that VCR adaptively modulates the magnitudes of gradients for updating the parameters of view-specific feature extractors to achieve a balanced multi-view learning procedure. In such a manner, BMvC achieves a better trade-off between the exploitation of view-specific patterns and the exploration of view-invariance patterns to fully learn the multi-view information for the clustering task. Finally, a set of experiments are conducted to verify the superiority of the proposed method compared with state-of-the-art approaches on eight benchmark MvC datasets. The demo code of this work is publicly available at https://github.com/guanyuezhen/BMvC.
@@ 1 Introduction
  • We perform extensive experiments on eight benchmark multi-view datasets to verify the effectiveness of the proposed method.
@@ 5 Conclusions
In this paper, we proposed a novel balanced multi-view clustering method to achieve more balanced multi-view learning and further improve the clustering performance. We first analyzed the imbalanced multi-view clustering problem existing in the joint-training paradigm from the gradient view. Then, we introduced a view-specific contrastive regularization to make a better trade-off between the exploitation of view-specific patterns and the exploration of view-invariance patterns to fully learn the multi-view information for the clustering task. Additionally, the theoretical analysis was provided to verify the effectiveness of the VCR from the gradient perspective. Finally, Extensive experiments on various benchmark multi-view clustering datasets were conducted to verify the efficacy of our method.